A metal alloy is 35% copper. Another metal alloy is 50% copper. How much of each alloy should be used to make 600 grams of a metal alloy that is 45% copper?
Found 3 solutions by jorel1380, ankor@dixie-net.com, ikleyn:
Answer by jorel1380(3719)

Let m be your first alloy, and n be your second. Then:
m+n=600
n=600-m
and
.35m+.5(600-m)=.45(600)
.35m+300-.5m=270
.15m=30
m=200 grams of the first alloy
600-m=400 grams of the second alloy. ☺☺☺☺

Answer by ankor@dixie-net.com(22740)

A metal alloy is 35% copper.
Another metal alloy is 50% copper.
How much of each alloy should be used to make 600 grams of a metal alloy that is 45% copper?
:
let x = amt of 50% alloy
resulting amt is be 600 gr, therefore
(600-x) = amt of 35% alloy
:
Amt of copper equation
.50x + .35(600-x) = .45(600)
.50x + 210 - .35x = 270
.50x - .35x = 270 - 210
.15x = 60
x = 60/.15
x = 400 gr of 50% alloy required
then, obviously;
600-400 = 200 gr of 35% alloy
:
;
Check this in the original equation
.50(400) + .35(200) = .45(600)
200 + 70 = 270

Answer by ikleyn(52785)

.
There is entire bunch of lessons covering various types of mixture problems
    - Mixture problems
    - More Mixture problems
    - Solving typical word problems on mixtures for solutions
    - Word problems on mixtures for antifreeze solutions
    - Word problems on mixtures for alloys (*)
    - Typical word problems on mixtures from the archive
in this site.

Read them and become an expert in solution the mixture word problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this textbook in the section "Word problems" under the topic "Mixture problems".